### v2

# Containers with non-uniform co-degree conditions, and an application to nearly orthogonal sets

Abstract
We develop a variant of the hypergraph container lemma with non-uniform conditions on the co-degrees. In particular, an upper bound on the co-degree of some subset of vertices $T$ is allowed to depend on where the vertices in $T$ live in the hypergraph, rather than having one condition which holds for all subsets. We use this to extend recent results on nearly-orthogonal sets in $\mathbb{F}_{p}^{d}$.
For a field $\mathbb{F}$ and integers $d,k$ and $\ell$, a set $A\subseteq\mathbb{F}^{d}$ is called $(k,\ell)$-nearly orthogonal if all vectors in $A$ are non-self-orthogonal and every $k+1$ vectors in $A$ contain $\ell+1$ pairwise orthogonal vectors. Recently, Haviv, Mattheus, Milojević and Wigderson have improved the lower bound on nearly orthogonal sets over finite fields, using counting arguments and a hypergraph container lemma. They showed that for every prime $p$ and an integer $\ell$, there is a constant $\delta(p,\ell)$ such that for every field $\mathbb{F}$ of characteristic $p$ and for all integers $d\geq k\geq\ell+1$, $\mathbb{F}^{d}$ contains a $(k,\ell)$-nearly orthogonal set of size $d^{\delta k/\log k}$. This nearly matches an upper bound $\binom{d+k}{k}$ coming from Ramsey theory. Moreover, they proved the same lower bound for the size of a largest set $A$ where for any two subsets of $A$ of size $k+1$ each, there is a vector in one of the subsets orthogonal to a vector in the other one. We prove a common generalisation of this result, showing that essentially the same lower bound holds for the size of a largest set $A\subseteq\mathbb{F}^{d}$ with the stronger property that given any family of subsets $A_{1},\ldots,A_{\ell+1}\subseteq A$, each of size $k+1$, we can find a vector in each $A_{i}$ such that they are all pairwise orthogonal.
1 Introduction
Consider an arbitrary field $\mathbb{F}$ and integers $d,k,\ell$ such that $k\geq\ell$. A set $A\subseteq\mathbb{F}^{d}$ of non-self-orthogonal vectors such that for any $k+1$ vectors there is a subset of size $\ell+1$ of pairwise orthogonal vectors is called a $(k,\ell)$-nearly orthogonal set. The largest size of such a set is denoted by $\alpha(d,k,\ell,\mathbb{F})$. The question of determining this number for $\mathbb{F}=\mathbb{R}$ was posed by Erdős in 1988 (see [10]).
An easy lower bound when $\ell=1$ can be found by considering the vectors of $k$ pairwise disjoint orthogonal bases of $\mathbb{R}^{d}$, showing that the minimal size is at least $kd$. Erdős conjectured this to be tight for $k=2$, which was proven in 1991 by Rosenfeld [21]. However, this lower bound is not tight for larger $k$, as was shown by Füredi and Stanley [15], proving the existence of a (5,1)-nearly-orthogonal set in $\mathbb{R}^{4}$ of size 24. A more general lower bound was given by Alon and Szegedy [4] (extending a result of Frankl and Rödl [14]), showing that for every integer $\ell$, there exists a constant $\delta(\ell)>0$ such that whenever $k\geq\max\{\ell,3\}$, $\alpha(d,k,\ell,\mathbb{R})\geq d^{\delta\log k/\log\log k}$.
More recently, Balla [5] considered a bipartite version of this problem when $\ell=1$. For integers $d,k$ and a field $\mathbb{F}$, let $\beta(d,k,\mathbb{F})$ denote the maximal size of a set $A\subseteq\mathbb{F}^{d}$ of non-self-orthogonal vectors such that for any two subsets $A_{1},A_{2}\subseteq A$ of size $k+1$ each, there exist $v_{1}\in A_{1}$ and $v_{2}\in A_{2}$ that are orthogonal. It clearly follows that $\alpha(d,k,1,\mathbb{F})\geq\beta(d,k,\mathbb{F})$. For this extended notion Balla proved that for all integers $d$ and $k\geq 3$, there exists a constant $\delta\geq 0$ such that $\beta(d,k,\mathbb{R})\geq d^{\delta\log k/\log\log k}$, matching the lower bound on $\alpha(d,k,1,\mathbb{R)}$. Nevertheless, this lower bound is still rather far away from the best current upper bound $\alpha(d,k,1,\mathbb{R})\leq O(d^{(k+1)/3)})$, proven by Balla, Letzter and Sudakov [6].
Besides $\mathbb{R}$, the maximal size of nearly orthogonal sets has also been studied in finite fields. This was first motivated by Codenotti, Pudlák and Resta [11], who showed the link between nearly-orthogonal sets and computational complexity of arithmetic circuits. Golovnev and Haviv [16] proved the existence of a constant $\delta>0$ such that $\alpha(d,2,1,\mathbb{F}_{2})\geq d^{1+\delta}$, showing a clear difference between nearly orthogonal sets in finite fields compared to the reals. This result was improved by Bamberg, Bishnoi, Ihringer and Ravi [9], giving an explicit construction of a (2,1)-nearly orthogonal set in $\mathbb{F}_{2}^{d}$ of size $d^{1.2895}$. Chawin and Haviv [10] considered the bipartite version and proved that for every prime $p$, there exists a positive constant $\delta(p)$ such that for every field $\mathbb{F}$ of characteristic $p$ and for all integers $k\geq 2$ and $d\geq k^{1/(p-1)}$, it holds that $\beta(d,k,\mathbb{F})\geq d^{\delta k^{1/(p-1)}/\log k}$. Hence, for $p=2$, it follows that
$\alpha(d,k,1,\mathbb{F}_{2})\geq\beta(d,k,\mathbb{F}_{2})\geq d^{\Omega(k/\log k)}.$ (1)
This almost matches an upper bound that follows easily from Ramsey theory, up to the logarithmic term in the exponent: consider an arbitrary $(k,1)$-almost orthogonal set in $\mathbb{F}^{d}$ and construct its orthogonality graph, where the vertices correspond to the vectors and there is an edge if the vectors are orthogonal. Clearly this graph does not contain a clique of size $d+1$, nor does it have an independent set of size $k+1$, because that would contradict that the set is $(k,1)$-almost orthogonal. Hence, the size of the set is upper bounded by the Ramsey number $R(d+1,k+1)$. The upper bound on $R(d+1,k+1)$ by Erdős and Szekeres [13] gives $\alpha(d,k,1,\mathbb{F}_{2})<\binom{d+k}{k}$. For $d$ sufficiently larger than $k$, this implies $\alpha(d,k,1,\mathbb{F}_{2})\leq d^{\Theta(k)}$.
In a recent paper, Haviv, Mattheus, Milojević and Wigderson [17] obtained the bound matching (1) for fields of characterstic larger than 2, and also extended it to the non-bipartite setup for arbitrary $\ell>1$.
Theorem 1 ([17]).
For every prime $p$ and integer $\ell\geq 1$, there exist constants $\delta_{1}(p),\delta_{2}(p,\ell)>0$, such that for every field $\mathbb{F}$ of characteristic $p$ and for all integers $d\geq k\geq\ell+1$, it holds that
$\beta(d,k,\mathbb{F})$ $\geq d^{\delta_{1}k/\log k}$
$\alpha(d,k,\ell,\mathbb{F})$ $\geq d^{\delta_{2}k/\log k}$
The proof of Theorem 1 is based on the probabilistic method. It relies on the fact that the orthogonality graph $G$ arising from finite fields has strong pseudorandom properties, and uses two different counting results:
  1. 1.
    Using an argument based on ideas of Alon and Rödl [2], Haviv et al. derived an upper bound on the number of pairs of sets in pseudorandom graphs with no edge across (used to bound $\beta$).
  2. 2.
    Using hypergraph containers [7, 22], they upper bounded the number of sets of size $k$ without a copy of $K_{\ell}$ (used to bound $\alpha$).
Having these counting results at hand, the proof proceeds by showing that a random induced subgraph of $G$ avoids sets of certain size which do not contain a copy of $K_{\ell}$, and it avoids pairs of subsets without an edge across. To make this actually work, a tensor-product trick along the lines of Alon and Szegedy [4] is applied.
1.1 Multipartite nearly-orthogonal sets
We unify the two above counting results by developing a container-type result for $\ell$-tuples of subsets without a crossing $K_{\ell}$ (Lemma 4). In particular, rather than counting the number of $\ell$-tuples of $k$-subsets without a crossing copy of $K_{\ell}$, we construct a small family $\mathcal{C}$ of small subsets with the property that for any ‘bad’ $\ell$-tuple $(U_{1},\ldots,U_{\ell})$, there exists $S\in\mathcal{C}$ such that $U_{i}\subseteq S$ for some $i\in[\ell]$.
The proof of Lemma 4 utilises a general hypergraph container result (Lemma 3) that allows for varying co-degree conditions. This is our main technical contribution. We believe that Lemma 3 is of independent interest, and it would be interesting to find further applications of this container lemma.
The application that we present in section 3 is an extension of the bounds on nearly orthogonal sets. Instead of considering $\alpha(d,k,\ell,\mathbb{F})$ and $\beta(d,k,\mathbb{F})$ separately, we consider a common generalisation $\beta(d,k,\ell,\mathbb{F})$ defined as the maximal size of a set $A\subseteq\mathbb{F}$ of non-self-orthogonal vectors such that for any $\ell+1$ subsets $A_{1},\dots,A_{\ell+1}\subseteq A$ of size $k+1$ each, there exist $v_{1}\in A_{1},\dots,v_{\ell+1}\in A_{\ell+1}$ that are all pairwise orthogonal. We prove bounds on $\beta(d,k,\ell,\mathbb{F})$ analogous to the ones in Theorem 1.
Theorem 2.
For every prime $p$ and integer $\ell\geq 1$, there exists a constant $\delta=\delta(p,\ell)>0$, such that for every field $\mathbb{F}$ of characteristic $p$ and for all integers $d\geq k\geq\ell+1$, it holds that
$\beta(d,k,\ell,\mathbb{F})\geq d^{\delta k/\log k}.$
Note that Theorem 2 implies Theorem 1. By taking $\ell=1$, we retrieve the bound on $\beta(d,k,\mathbb{F})$ and by setting $A_{1}=\dots=A_{\ell+1}$, we retrieve the bound on $\alpha(d,k,\ell,\mathbb{F})$.
2 A container lemma with varying co-degree condition
Hypergraph containers, developed by Balogh, Morris and Samotij [7] and Saxton and Thomason [22], provide a framework for studying independent sets in hypergraphs. On a high level, the main idea of hypergraph containers is to show that given any independent set $I$ in a hypergraph $\mathcal{H}$, one can carefully choose a small subset $F\subseteq I$, called a fingerprint, such that just based on $F$ one can confine $I$ to some $C\subseteq V(\mathcal{H})$, called a container. Importantly, the size of $C$ is bounded away from $|V(\mathcal{H})|$. This innocent looking statement has far reaching consequences. For a thorough introduction, see [8].
To make use of the standard hypergraph container framework, one typically needs that the given hypergraph has fairly uniform degree and co-degree distribution. This will not be the case in our application. In particular, our hypergraph is $\ell$-partite and degrees depend on which part the vertex belongs to. More generally, the number of edges sitting on a given set of vertices depends on which parts these vertices lie in. Lemma 3 is stated with this setup in mind.
The following statement is inspired by the one used by Nenadov and Pham [20]. Let $V$ be an arbitrary set. For a subset $T\subseteq V$, we let $\langle T\rangle$ denote the upset of $T$, i.e. $\langle T\rangle=\{S\subseteq V:T\subseteq S\}$.
Lemma 3.
Let $\ell\in\mathbb{N}$ and let $\mathcal{H}$ be an $\ell$-partite $\ell$-uniform hypergraph on the vertex set $(V_{1},\dots,V_{\ell})$. Suppose there exist $p_{1},\dots,p_{\ell}\in(0,1]$, $K>1$ and a probability measure $\nu$ over $2^{V}$, $V=\bigcup_{i}V_{i}$, supported on $E(\mathcal{H})$ such that for all sets $S\subseteq[\ell]$, $T\in\prod_{i\in S}V_{i}$ and $i\in S$, we have
$\nu(\langle T\rangle)\leq\frac{K}{|V_{i}|}\prod_{j\in S\setminus\{i\}}p_{j}.$ (2)
Let $U=(U_{1},\dots,U_{\ell})$, $U_{i}\subseteq V_{i}$, be an arbitrary independent set in $\mathcal{H}$ such that $|U_{j}|\geq p_{j}|V_{j}|$ for all $j\in[\ell]$. Then,
  1. (i)
    there is a unique $\ell$-tuple $\mathbf{F}=(\emptyset,F_{2},\ldots,F_{\ell})$ associated with $(U_{1},\ldots,U_{\ell})$, where $F_{j}\subseteq U_{j}$ and $|F_{j}|=|V_{j}|p_{j}$ for $2\leq j\leq\ell$;
  2. (ii)
    there is an index $i=i(\mathbf{F})\in[\ell]$, a corresponding set $C_{i}=C_{i}(\mathbf{F})\subseteq V_{i}$ and a constant $\zeta=\zeta(K,\ell)>0$ such that $|C_{i}|\leq(1-\zeta)|V_{i}|$ and $U_{i}\subseteq F_{i}\cup C_{i}$.
  3. (iii)
    Furthermore, given any $U_{1}^{\prime},\ldots,U_{\ell}^{\prime}$ such that $F_{i}\subseteq U_{i}^{\prime}\subseteq U_{i}$, the $\ell$-tuple $\mathbf{F}=(\emptyset,F_{2},\ldots,F_{\ell})$ associated with $(U_{1},\ldots,U_{\ell})$ is also associated with $(U_{1}^{\prime},\ldots,U_{\ell}^{\prime})$.
Whenever $\nu$ is the uniform probability measure over $\mathcal{H}$, i.e. $\nu(X)=1/e(\mathcal{H})$ if $X\in E(\mathcal{H})$ and $\nu(X)=0$ otherwise, this condition is just a co-degree condition analogous to the one in [7, 22]. Indeed, we have
$\nu(\langle T\rangle)=\sum_{S\subseteq V:T\subseteq S}\nu(S)=\Delta_{T}\cdot\frac{1}{e(\mathcal{H})},$
where $\Delta_{T}$ denotes the codegree of the set $T$, that is, the number of hyperedges in $\mathcal{H}$ which contain $T$. As remarked earlier, the main distinction between Lemma 3 and its classical predecessors is that we allow for more freedom of co-degrees of different sets. Usually, one considers a single condition along the lines of (2) that should hold for all sets $T$, independent of where $T$ lives in the hypergraph. In Lemma 3, the spread condition depends on where $T$ lives, allowing for more variation in this measure between different sets. This possibility for more variation is necessary for our application to nearly-orthogonal sets.
Note that Lemma 3 bears some resemblance to the asymmetric container lemma by Morris, Samotij and Saxton [19]. Similarly to the statement in [19], it is straightforward to extend the proof of Lemma 3 to support hypergraphs which are not necessarily $\ell$-uniform but where hyperedges may contain more than a single vertex in each of the sets $V_{i}$. As we do not have an application for such a statement, we have decided to state a simpler version that is sufficient for us, which is the proof Lemma 4. That being said, the hypergraph that arises in the proof of Lemma 4 is very dense and allows one to take $p_{i}=1/|V_{i}|$, which amounts to each $F_{i}$ being a single vertex. Proving a version of Lemma 3 for this special case would be a bit simpler, but as we hope that it will find further applications, we gave a more general version. The proof very closely follows a recent proof of the original hypergraph containers by Nenadov and Pham [20].
Proof of Lemma 3.
We prove the lemma by induction on $\ell$. To that end, first consider $\ell=1$. Let $C_{1}\subseteq V_{1}$ to be the set of all vertices $v\in V_{1}$ with $\nu(\{v\})=0$. Since $\nu$ is supported on $E(\mathcal{H})$, any independent set avoids $V_{1}\setminus C_{1}$. There are at least $|V_{1}|/K$ vertices with non-zero measure, so the lemma holds for $\zeta=1/K$. We proceed to the case where $\ell\geq 2$. We use the following notation: given a subset $X\subseteq V$, let $\mathcal{I}_{\mathcal{H}}(X)$ denote the set of all indices for which $V_{i}$ intersects $X$, that is,
$\mathcal{I}_{\mathcal{H}}(X):=\{i\in[\ell]:X\cap V_{i}\neq\emptyset\}.$
Constructing fingerprint $F_{\ell}$.
Set $F_{\ell}=\emptyset\subseteq U_{\ell}$, $\mathcal{L}=\emptyset\subseteq 2^{V\setminus V_{\ell}}$, $\mathcal{D},\mathcal{H}^{\prime}=\emptyset\subseteq\mathcal{H}$ and $\mathcal{R}=\mathcal{H}$. Repeat the following steps $|V_{\ell}|p_{\ell}$ times:
  • Pick the vertex $v\in U_{\ell}\setminus F_{\ell}$ with the largest value $\nu(\langle v\rangle\cap\mathcal{R})$. Tie-breaking is done in some canonical way, that is, following an ordering of the vertices agreed upon at the beginning of the algorithm. Note that we can always pick such a vertex $v$ since $|U_{\ell}|\geq p_{\ell}|V_{\ell}|$.
  • Add $v$ to $F_{\ell}$ and set $\mathcal{H}^{\prime}=\mathcal{H}^{\prime}\cup(\langle v\rangle\cap\mathcal{R})$.
  • For each $X\in 2^{V\setminus V_{\ell}}\setminus\mathcal{L}$ of size $|X|\leq\ell-1$, such that there exists $s\in\mathcal{I}_{\mathcal{H}}(X)$ for which the following inequality holds:
    $\nu(\langle X\rangle\cap\mathcal{H}^{\prime})>\frac{K}{|V_{s}|}\cdot p_{\ell}\prod_{i\in\mathcal{I}_{\mathcal{H}}(X)\setminus\{s\}}p_{i},$ (3)
    add $X$ to $\mathcal{L}$, set $\mathcal{D}=\mathcal{D}\cup(\langle X\rangle\cap\mathcal{R})$ and set $\mathcal{R}=\mathcal{H}\left[V\setminus F_{\ell}\right]\setminus\mathcal{D}$.
Observations on the process.
Before we show how to choose other fingerprints $F_{2},\ldots,F_{\ell-1}$ and construct a container $C_{i}$ for some $U_{i}$, we first make some observations about the described process. First of all, for any subset $X\subseteq V\setminus V_{\ell}$, because of (2), the value of $\nu(\langle X\rangle\cap\mathcal{H}^{\prime})$ increases by at most
$\nu(\langle X\cup v\rangle)\leq\frac{K}{|V_{\ell}|}\cdot p_{\ell}\prod_{i\in\mathcal{I}_{\mathcal{H}}(X)}p_{i}$
after adding $v$ to $F_{\ell}$. Note that the value of $\nu(\langle X\rangle\cap\mathcal{H}^{\prime})$ does not change anymore once $X$ is added to $\mathcal{L}$, and this happens when (3) is satisfied. Therefore, at the end of the process it holds that for every $X\subseteq V\setminus V_{\ell}$ of size $|X|\leq\ell-1$, including those that are in $\mathcal{L}$, and for all $s\in\mathcal{I}_{\mathcal{H}}(X)$, we have
$\nu(\langle X\rangle\cap\mathcal{H}^{\prime})\leq\frac{2K}{|V_{s}|}\cdot p_{\ell}\prod_{i\in\mathcal{I}_{\mathcal{H}}(X)\setminus\{s\}}p_{i}.$ (4)
Secondly, given $U_{\ell}^{\prime}\subseteq U_{\ell}$ such that $F_{\ell}\subseteq U_{\ell}$, if we start the process with $U^{\prime}_{\ell}$ instead of $U_{\ell}$, it produces the same output $F_{\ell}$, as well as the same $\mathcal{H}^{\prime}$ and $\mathcal{R}$.
Next, we derive a lower bound on $\nu(\mathcal{H}^{\prime})$. To start, we know that for each edge $e\in\mathcal{D}$, there exists $X\in\mathcal{L}$ such that $e\in\langle X\rangle$. Thus,
$\sum_{X\in\mathcal{L}}\nu(\langle X\rangle)\geq\nu(\mathcal{D}).$
Also, we know by (3), that for each $X\in\mathcal{L}$ there is an $s(X)\in\mathcal{I}_{\mathcal{H}}(X)$ such that
[…]
Moreover, each edge $e\in\mathcal{H}^{\prime}$ contributes at most to $2^{\ell}$ terms in $\nu(\langle X\rangle\cap\mathcal{H}^{\prime})$. So,
$\nu(\mathcal{H}^{\prime})\geq 2^{-\ell}\sum_{X\in\mathcal{L}}\nu(\langle X\rangle\cap\mathcal{H}^{\prime})>2^{-\ell}p_{\ell}\nu(\mathcal{D}).$
In conclusion, we have now derived a relationship between $\nu(\mathcal{H}^{\prime})$ and $\nu(\mathcal{D})$:
$\nu(\mathcal{H}^{\prime})>2^{-\ell}p_{\ell}\nu(\mathcal{D}).$ (5)
Secondly, we relate the measure on $\mathcal{H}^{\prime}$ to the measure on the vertices in $U_{\ell}$ that are not contained in the fingerprint $F_{\ell}$. Let $\mathcal{R}_{i}$ denote the hypergraph $\mathcal{R}$ at the moment when the $i$-th vertex $v_{i}$ is added to $F_{\ell}$. Since $\mathcal{R}$ is non-increasing and we always chose $v_{i}$ with maximal value $\nu(\langle v\rangle\cap\mathcal{R})$, we have
$\nu(\mathcal{H}^{\prime})=\sum_{i=1}^{|F_{\ell}|}\nu(\langle v_{i}\rangle\cap\mathcal{R}_{i})\geq\sum_{i=1}^{|F_{\ell}|}\max_{v\in U_{\ell}\setminus F_{\ell}}\nu(\langle v\rangle\cap\mathcal{R}).$
Therefore, we can conclude that
$\nu(\mathcal{H}^{\prime})\geq|V_{\ell}|p_{\ell}\max_{v\in U_{\ell}\setminus F_{\ell}}\nu(\langle v\rangle\cap\mathcal{R}).$ (6)
Constructing the container.
Set $\alpha=2^{-\ell-2}$ and consider two cases.
Case 1: $\nu(\mathcal{H}^{\prime})<\alpha p_{\ell}$. By (5), we have $\nu(\mathcal{D})<\frac{1}{4}$, so, using $\mathcal{H}=\mathcal{H}^{\prime}\cup\mathcal{R}\cup\mathcal{D}$, $\nu(\mathcal{R})>\frac{1}{2}$. By (6) it holds that for every $v\in U_{\ell}\setminus F_{\ell}$,
$\nu(\langle v\rangle\cap\mathcal{R})\leq\frac{\alpha}{|V_{\ell}|}.$
We use this measure to define our container. Let $C_{\ell}$ be the set of all vertices $v\in V_{\ell}\setminus F_{\ell}$ such that $\nu(\langle v\rangle\cap\mathcal{R})\leq\frac{\alpha}{|V_{\ell}|}$. Clearly, we have $U_{\ell}\setminus F_{\ell}\subseteq C_{\ell}$. Moreover, by (2),
$\nu(\mathcal{R})\leq\sum_{v\in C_{\ell}}\nu(\langle v\rangle\cap\mathcal{R})+\sum_{w\in V_{\ell}\setminus(F_{\ell}\cup C_{\ell})}\nu(\langle w\rangle\cap\mathcal{R})<\alpha+(|V_{\ell}|-|C_{\ell}|)\frac{K}{|V_{\ell}|}.$
Hence,
$|C_{\ell}|<|V_{\ell}|-(\nu(\mathcal{R})-\alpha)\frac{|V_{\ell}|}{K}<(1-\zeta)|V_{\ell}|$
for $\zeta=1/(4K)$. Take $F_{j}\subseteq U_{j}$ to be a subset of size $|V_{j}|p_{j}$ chosen in some canonical way, for $2\leq j<\ell$. These fingerprints are defined for the convenience of having all fingerprints of the specified size. So we have obtained the $\ell$-tuple of fingerprints $\mathbf{F}=(\emptyset,F_{2},\dots,F_{\ell})$ and index $i(\mathbf{F})=\ell$ such that all conditions of the lemma are satisfied.
Case 2: $\nu(\mathcal{H}^{\prime})\geq\alpha p_{\ell}$. We use the induction hypothesis to find fingerprints for $U_{1},\dots,U_{\ell-1}$ and a container for one of these sets. Let $\mathcal{H}^{\prime\prime}$ be the $(\ell-1)$-uniform $(\ell-1)$-partite hypergraph consisting of sets $X$ such that $X=H^{\prime}\setminus F_{\ell}$ for some $H^{\prime}\in\mathcal{H}^{\prime}$. Let $\nu^{\prime}$ be the probability measure over $2^{V\setminus V_{\ell}}$ given by
$\nu^{\prime}(X)\propto\left\{\begin{array}[]{ll}\nu((X\cup 2^{F_{\ell}})\cap\mathcal{H}^{\prime})&\mbox{if }X\in\mathcal{H}^{\prime\prime}\\ 0&\mbox{otherwise.}\end{array}\right.$
Consider some $X\in\mathcal{H}^{\prime\prime}$. For $s\in\mathcal{I}_{\mathcal{H}^{\prime\prime}}(X)$, we have
[…]
for $K^{\prime}=2K/\alpha$. Furthermore, $(U_{1},\dots,U_{\ell-1})$ is an independent set in $\mathcal{H}^{\prime\prime}$. Thus, by the induction hypothesis, there exists $(F_{2},\ldots,F_{\ell-1})$ such that $F_{j}\subseteq U_{j}$ and $|F_{j}|=|V_{j}|p_{j}$ for all $12\leq j\leq\ell-$ 1, and a suitable container $C_{i}=C_{i}(F_{2},\dots,F_{\ell-1})$ for some $i\in[\ell-1]$ such that $U_{i}\subseteq C_{i}\cup F_{i}$ and $|C_{i}|\leq(1-\zeta)|V_{i}|$.
3 Nearly Orthogonal Sets
The following result has the role of Theorem 2.4 and Theorem 2.6 in [17]. Unlike these two theorems, here we do not count the number of ‘bad’ tuples of sets. Rather, we give an upper bound on the size of a family which underpins such tuples. To this end, we will need the notion of $(n,D,\lambda)$-graphs. Recall that a $D$-regular graph on $n$-vertices is an $(n,D,\lambda)$-graph if the absolute values of all but the largest eigenvalue of its adjacency matrix are at most $\lambda$. Given a graph $G$, we denote with $\mathfrak{I}_{\ell}(G)$ the family of all $\ell$-tuples $(U_{1},\ldots,U_{\ell})$, where $U_{i}\subseteq V(G)$ for each $i\in[\ell]$, such that there is no copy $(v_{1},\ldots,v_{\ell})$ of $K_{\ell}$ in $G$ with $v_{i}\in U_{i}$ for $i\in[\ell]$.
Lemma 4.
Let $c\in(0,1/2]$ be a constant and $n$ a sufficiently large integer. Suppose $G$ is an $(n,D,\lambda)$-graph with $D\geq cn$. There exists $C=C(\ell,c)>0$ and a family $\mathcal{C}$ of subsets of $V(G)$ with the following properties:
  • $|\mathcal{C}|\leq n^{C\log n}$;
  • each $S\in\mathcal{C}$ is of size $|S|\leq C\lambda$;
  • for every $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$, there exists $S\in\mathcal{C}$ and $i\in[\ell]$ such that $U_{i}\subseteq S$.
Note that we make no assumptions on $\lambda$ in Lemma 4. However, as $D\geq cn$, for $\lambda>|D|/(c\cdot C)$ we have $C\lambda>n$, thus the upper bound on $S\in\mathcal{C}$ becomes trivial.
Proof of Lemma 4.
Form a hypergraph $\mathcal{H}$ by taking $\ell$ disjoint copies of $V(G)$, named $V_{1},\ldots,V_{\ell}$, and putting a hyperedge on top of $(v_{1},\ldots,v_{\ell})$, $v_{i}\in V_{i}$ for $i\in[\ell]$, if the corresponding vertices in $G$ form a copy of $K_{\ell}$.
Using the Expander Mixing Lemma [3, Corollary 9.2.5] and a standard embedding argument [^3] See e.g.  [18, Theorem 4.10] or [12, Proposition 3.3]. Note that these results show both a lower and upper bound. We only require the lower bound, which can be more easily obtained with an inductive argument., for any $\ell$-tuple of subsets $(W_{1},\dots,W_{\ell})$, $W_{i}\subseteq V_{i}$, we have
$e(\mathcal{H}[W_{1}\cup\ldots\cup W_{\ell}])\geq\frac{1}{2}\prod_{i=1}^{\ell}|W_{i}|\left(\frac{D}{n}\right)^{\binom{\ell}{2}}>c^{\ell^{2}}\prod_{i=1}^{\ell}|W_{i}|,$ (7)
provided $|W_{i}|\geq C^{\prime}\lambda\left(n/D\right)^{\ell-1}$ for large enough constant $C^{\prime}$.
Note that if $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$ then $(U_{1},\ldots,U_{\ell})$ is an independent set in $\mathcal{H}$. First of all, add to $\mathcal{C}$ all subsets of $V(G)$ of size $C\log n$, where $C=C^{\prime}c^{-\ell}\ell^{2}\zeta^{-1}$. The meaning of the different factors in this constant will become apparent in different later steps. This takes care of all $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$ with some $|U_{i}|\leq C\log n$. Consider now some $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$ with $|U_{i}|>C\log n$ for all $i\in[\ell]$. Set $W_{i}=V_{i}$, $U_{i}^{\prime}=U_{i}$ and $F_{i}=\emptyset$ for $i\in[\ell]$, and as long as each $W_{i}$ is of size at least $C\lambda\left(n/D\right)^{\ell-1}$, repeat the following:
  • Let $\mathcal{H}^{\prime}=\mathcal{H}[W_{1}\cup\ldots\cup W_{\ell}]$, and define a measure $\nu$ on $2^{V(\mathcal{H}^{\prime})}$ to be uniform on $\mathcal{H}^{\prime}$:
    $\nu(X)=\left\{\begin{array}[]{ll}\frac{1}{e(\mathcal{H}^{\prime})}&\text{if }X\in\mathcal{H}^{\prime}\\ 0&\text{otherwise.}\end{array}\right.$
    By (7), the condition (2) is satisfied for $K=c^{-\ell^{2}}$ and $p_{i}=1/|W_{i}|$ for each $i\in[\ell]$.
  • As $(U_{1}^{\prime},\ldots,U^{\prime}_{\ell})$ is an independent set in $\mathcal{H}^{\prime}$, apply Lemma 3 to obtain fingerprints $(\emptyset,F^{\prime}_{2}\ldots,F^{\prime}_{\ell})$. Let $i\in[\ell]$ be the index and $C_{i}\subseteq W_{i}$ the set corresponding to $(\emptyset,F^{\prime}_{2},\ldots,F^{\prime}_{\ell})$. Note that for each $j\in[2,\ell]$, $F^{\prime}_{j}$ is a single vertex, and $|C_{i}|\leq(1-\zeta)|W_{i}|$, where $\zeta=\zeta(K,\ell)>0$. Moreover, $U_{i}^{\prime}\subseteq F^{\prime}_{i}\cup C_{i}$.
  • Set $U_{i}^{\prime}=U_{i}^{\prime}\cap C_{i}$, $W_{i}=C_{i}$, and $F_{j}:=F_{j}\cup F^{\prime}_{j}$ for $j\in[2,\ell]$. Proceed to the next round.
As the sets $W_{i}$ are shrinking, the hypergraph $\mathcal{H}^{\prime}$ might become increasingly more unbalanced. The average degree of a vertex in $W_{i}$ is $\Theta(\prod_{j\neq i}|W_{j}|)$, meaning that if $|W_{i}|\ll|W_{i^{\prime}}|$, then the average degree of vertices in $W_{i}$ is significantly larger than of those in $W_{i^{\prime}}$. This dependency is captured by the parameter $p_{i}$ and is the main reason the standard hypergraph containers do not apply.
As one of the $W_{i}$’s shrink by a factor of at least $1-\zeta$ in each round, the whole process terminates after at most $t=\ell\zeta^{-1}\log n$ iterations. The size of each $F_{i}$ is thus at most $t$, and for convenience we add to each $F_{i}$ some elements from $U_{i}\setminus F_{i}$, in a canonical way, so that $|F_{i}|=t$. Let $z\in[\ell]$ denote the index of the set $W_{z}$ which is smaller than $C^{\prime}\lambda(n/D)^{\ell-1}$ at the end of the process. By the property iii of Lemma 3, we can reproduce each iteration of the process if we start with $(\emptyset,F_{2},\ldots,F_{\ell})$ instead of $(U_{1},\ldots,U_{\ell})$ (it is interesting to note that $U_{1}$ plays no role in the process). Therefore, $z$ and $W_{z}$ are indeed a function of $(F_{2},\ldots,F_{\ell})$. Finally, note that for each $j\in[\ell]$, $U_{j}\subseteq F_{j}\cup W_{j}$ (where $F_{j}=\emptyset$ for $j=1$) and so specifically this is also true for $j=z$. When $n$ is large enough so that $\lambda\gg t$, we obtain
$|F_{z}\cup W_{z}|\leq t+C^{\prime}c^{-\ell}\lambda<C\lambda.$
We expand the family of subsets $\mathcal{C}$ in the following way: for each $\mathbf{F}=(\emptyset,F_{2},\ldots,F_{\ell})$ such that $|F_{2}|=\ldots=|F_{\ell}|=t$ and $\mathbf{F}$ is the resulting fingerprint when running the process with some $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$, let $i=i(\mathbf{F})\in[\ell]$ and $W_{i}=W_{i}(\mathbf{F})\subseteq V(G)$, and add $F_{i}\cup W_{i}$ to $\mathcal{C}$. This guarantees that for every $(U_{1},\ldots,U_{\ell})\in\mathfrak{I}_{\ell}(G)$, one of the sets $U_{i}$ is a subset of some $S\in\mathcal{C}$. The size of $\mathcal{C}$ is at most
$\binom{n}{C\log n}+\binom{n}{\ell\zeta^{-1}\log n}^{\ell}<n^{C\log n}.$
With Lemma 4 at hand, the proof of Theorem 2 follows the structure of Haviv et al. [17].
Proof of 2..
Let $C=C(\ell,1/2p)>0$ be the constant given by Lemma 4, and choose $t$ to be the largest integer such that $C\log(p^{t})<k/6$. Furthermore, we assume $k$ is large enough such that we can apply Lemma 4 with $n=p^{t}$. For small $k$ we pick $\delta(p,\ell)$ to be sufficiently small to satisfy Theorem 2 when using $d$ basis vectors in $\mathbb{F}_{p}^{d}$. Let $m\geq 1$ be the largest integer such that $t^{m}\leq d$.
Let $Q$ denote the set of non-self-orthogonal distinct vectors in $\mathbb{F}_{p}^{t}$. Note that $|Q|\geq p^{t-1}$. Let $V\subseteq Q^{m}$ be a set of
$r=\lfloor p^{mt/(4\ell)}\rfloor$
$m$-tuples of vectors from $Q$, chosen uniformly at random and independently. Let us denote these $m$-tuples as $(v_{1}^{(i)},\ldots,v_{m}^{(i)})$, for $i\in[r]$. Note that it is possible that some of these $m$-tuples are identical.
Consider the orthogonality graph $G(p,t)$ which has vertex set equal to the vectors in $\mathbb{F}_{p}^{t}$ and $u\sim v$ if $\langle u,v\rangle=0$. It is well known that this graph is an $(n,D,\lambda)$-graph (e.g. see [1, Page 220]) with the following parameters:
$n=p^{t}-1,\hskip 10.0ptD=p^{t-1}-1,\hskip 5.0pt\text{and}\hskip 7.0pt\lambda=p^{t/2-1}(p-1).$
Since $t$ is sufficiently large, we can apply Lemma 4. Let $\mathcal{C}$ be the family given by Lemma 4 together with constant $C$. Recall that $|\mathcal{C}|\leq n^{C\log n}$ and that all elements in $\mathcal{C}$ have size at most $C\lambda$.
Given a set $K=\{i_{1},\ldots,i_{k}\}\subseteq[r]$ of size $|K|=k$ and $z\in[m]$, let us denote with $\mathcal{B}_{z}(K)$ the event that $\{v_{z}^{(i_{1})},\ldots,v_{z}^{(i_{k})}\}\subseteq S$ for some $S\in\mathcal{C}$. By the union bound, we have
$\Pr[\mathcal{B}_{z}(K)]$ $\leq\sum_{S\in\mathcal{C}}\Pr\left[\{v_{z}^{(i_{1})},\ldots,v_{z}^{(i_{k})}\}\subseteq S\right]$
$=\sum_{S\in\mathcal{C}}\left(\frac{|S|}{|Q|}\right)^{k}\leq|\mathcal{C}|\left(\frac{C\lambda}{p^{t-1}}\right)^{k}$
$\leq 2^{C\log^{2}(p^{t})}\left(Cp^{-t/2+1}\right)^{k}<\left(Cp^{-t/3+1}\right)^{k}.$
Let $\mathcal{B}(K)$ denote the event that $\mathcal{B}_{z}(K)$ happens for at least $m/\ell$ different values of $z\in[m]$. By union bound and the fact that the events $\mathcal{B}_{z}(K)$ are independent for different values of $z$, we have
$\Pr[\mathcal{B}(K)]\leq\sum_{\begin{subarray}{c}Z\subseteq[m]\\ |Z|\geq m/\ell\end{subarray}}\Pr\left[\bigwedge_{z\in Z}\mathcal{B}_{z}(K)\right]\leq 2^{m}\left(Cp^{-t/3+1}\right)^{km/\ell}\leq\left(2Cp^{-t/3+1}\right)^{km/\ell},$
where the last inequality follows from $k\geq\ell$. Finally, let $\mathcal{B}$ denote the event that $\mathcal{B}(K)$ happens for some $K\subseteq[r]$ of size $k$. By one last application of the union bound, we have
$\Pr[\mathcal{B}]\leq\sum_{\begin{subarray}{c}K\subseteq[r]\\ |K|=k\end{subarray}}\Pr[\mathcal{B}(K)]\leq\binom{r}{k}\left(2Cp^{-t/3+1}\right)^{km/\ell}\leq p^{kmt/(4\ell)}\left(2Cp^{-t/3+1}\right)^{km/\ell}\leq\left(2Cp^{-t/12+1}\right)^{km/\ell}<1.$
The last inequality holds for large enough $t$, which is implied by the assumption that $k$ is large.
To summarise, there is a choice for $V$ for which $\mathcal{B}$ does not occur. Now recall that for two vectors $u=(u_{1},\ldots,u_{t}),v=(v_{1},\ldots,v_{t})\in\mathbb{F}^{t}$, their tensor product $u\otimes v=w\in\mathbb{F}^{t^{2}}$ is the vector whose entries are indexed by $(a,b)\in[t]^{2}$ and are defined by $w_{(a,b)}=u_{a}v_{b}$. Form the set $W$ by taking the tensor product $w_{i}=v_{1}^{(i)}\otimes\ldots\otimes v_{m}^{(i)}$ for each $i\in[r]$. Note that all the vectors in $W$ are non-self-orthogonal, and if $\langle v_{z}^{(i)},v_{z}^{(j)}\rangle=0$ for some $z\in[m]$ and $i,j\in[r]$, then $\langle w_{i},w_{j}\rangle=0$. We claim that the corresponding set $W$ satisfies that every $\ell$-tuple of subsets $(W_{1},\dots,W_{\ell})$, with each $W_{i}$ of size $k$, contains $\ell$ vectors $\{w_{i}\in W_{i}\}_{i\in[\ell]}$ that are pairwise orthogonal. Consider one such $\ell$-tuple, and let $K_{i}\subseteq[r]$ denote the indices of vectors in $W_{i}$. As none of the events $\mathcal{B}(K_{i})$ happen, the set $Z_{i}\subseteq[m]$ of values $z\in[m]$ for which $\mathcal{B}_{z}(K_{i})$ happens is of size $|Z_{i}|<m/\ell$. Therefore, there exists $z\in[m]$ such that none of the events $\mathcal{B}_{z}(K_{1}),\ldots,\mathcal{B}_{z}(K_{\ell})$ happen. Let $U_{i}=\{v_{z}^{(j)}\colon j\in K_{i}\}$ for each $i\in[\ell]$. Then $(U_{1},\ldots,U_{\ell})\not\in\mathfrak{I}_{\ell}(G(p,t))$ as otherwise we would have $U_{i}\subseteq S$ for some $i\in[\ell]$ and $S\in\mathcal{C}$, i.e. $B_{z}(K_{i})$ happens for some $i$, which is not the case. This means there is $i_{j}\in K_{j}$ for each $j\in[\ell]$ such that all the vectors in $\{v_{z}^{(i_{1})},\ldots,v_{z}^{(i_{\ell})}\}$ are pairwise orthogonal. By the observed property of tensor products, we consequently have that all the vectors in $\{w_{i_{1}},\ldots,w_{i_{\ell}}\}$ are also pairwise orthogonal.
To conclude, we have obtained a set $W\subseteq\mathbb{F}_{p}^{mt}$ of size
$|W|=r=p^{mt/(4\ell)}=p^{\Theta(\log d\cdot t/\log t)}=d^{\Theta(k/\log k)}$
with the desired property. Now we add $d-t^{m}$ zeroes at the end of each vector in $W$ such that we obtain a subset of $\mathbb{F}_{p}^{d}$ and conclude the proof. ∎